OpenStudy (anonymous):

What would happen if CH4 was removed from the following reaction that has reached equilibrium? CO (g) + 3H2 (g) <--> CH4 (g) + H2O (g) a) The reaction would not be affected. b) The equilibrium would shift to the right. c) The equilibrium would shift to the left. d) The product concentration would equal the reactants.

OpenStudy (anonymous):

b \[ K_p = \frac{[CH_4] \cdot [H_2O]}{[CO] \cdot [H_2]^3} \] Kp is a constant. Now if you remove CH4, the numerator becomes less and since Kp is constant the denominator also has to become less i.e. CO & H2 has to be get consumed and hence the equilibrium will shift forward.

OpenStudy (anonymous):

Thank you so much!